An intense stream of water of cross section area A strikes a wall at an angle X with the normal to the wall return back elastically.if the density of water Is d and it’s velocity is v, then the force exerted in the wall will be
1
Expert's answer
2019-09-09T11:00:10-0400

Impulse on wall due to infinitesimal mass of the water is


"I=\Delta P=dm\ v\cos{X}-(-dm\ v\cos{X})=2\ dm\ v\cos{X}"

The force exerted in the wall will be


"F=\frac{\Delta P}{\Delta t}=2v\cos{X}\frac{d m}{d t}=2v\cos{X}\left(dA\frac{d x}{d t}\right)"

"F=2v\cos{X}(dAv)=2dAv^2\cos{X}"